Silicon dioxide with the chemical formula SiO2. how many moles of silicon dioxide are there in 25-3g of silicon dioxide


1
Expert's answer
2021-03-12T05:58:24-0500

1) 25^-3=0,000064 g SiO2

2)n(SiO2)=m(SiO2)/M(SiO2)= 0,000064/(60)=0,000001067 mole SiO2